OpenStudy (anonymous):

Evaluate: 4 − 49 ÷ 7 + 3^2

OpenStudy (anonymous):

49/7 =7 3^2 =3x3=9 4-7+9=-3+9=......

OpenStudy (anonymous):

so im suppose to be doing order of operations right

OpenStudy (anonymous):

yup multiplication and division first then, plus and minus

OpenStudy (anonymous):

4-7+9=-3+9= would it be 6

OpenStudy (anonymous):

yes, you got it!!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

how do you simpfly −2(2u − 1) using distributive property

OpenStudy (anonymous):

(-2)(2u)+(-1)(-2)=-4u+2

OpenStudy (anonymous):

how do you do that so fast

OpenStudy (anonymous):

can you help me simple this using distributive property−3(2d − 1) + 4d

OpenStudy (anonymous):

-3(2d -1)+4d =(-3)(2d) + (-3)(-1) +4d =-6d +3 +4d =.....

OpenStudy (anonymous):

i dont get how to do this at all

OpenStudy (anonymous):

do you know take out common factor such as ab+ac =a(b+c)?

OpenStudy (anonymous):

my problem is not understanding each step

OpenStudy (anonymous):

ab+ac = a(b+c) <-- this is taking out common factor for expansion, a(b+c) multiply a to each of the components in the bracket ab and ac so a(b+c)=ab+bc

OpenStudy (anonymous):

-3(2d -1) multiply (-3) to each of the components in the bracket (-3)(2d) = -6d and (-3)(-1) =3 -3(2d -1) = -6d+3